𝑦′ βˆ’ 𝑦 = 𝑒2π‘₯𝑦3


1
Expert's answer
2022-02-13T17:49:03-0500
"y'-y = e^{2x}y^3"

Substitute "v=y^{1-3}"


"v'=-2y^{-3}y'"

"-\dfrac{1}{2}v'-v=e^{2x}"

"v'+2v=-2e^{2x}"

Integrating factor


"\mu(x)=e^{\int(2)dx}=e^{2x}"

"e^{2x}v'+2e^{2x}v=-2e^{4x}"

"d(e^{2x}v)=-2e^{4x}dx"

Integrate


"\int d (e^{2x}v)=-\int2e^{4x}dx"

"e^{2x}v=-\dfrac{1}{2}e^{4x}+C"

"v=-\dfrac{1}{2}e^{2x}+\dfrac{1}{2}Ce^{-2x}"

Then


"\dfrac{1}{y^2}=-\dfrac{1}{2}e^{2x}+\dfrac{1}{2}Ce^{-2x}"


"y^2=\dfrac{2}{Ce^{-2x}-e^{2x}}"

"y^2=\dfrac{2e^{2x}}{C-e^{4x}}"